# Not all Layers of LLMs are Necessary during Inference

### v2

@@ 4.2 Classifier
The Classifier determines if the signal is compelling enough to warrant an early termination of the process. The rule-based approach heavily relies on rules, and the cost of individually constructing domain-specific features is high Huang et al. (2017); Yang et al. (2020); Wang et al. (2022). Conversely, the plug-and-play nature of the gating function Lin et al. (2017); Bejnordi et al. (2019) provides greater universality. Nonetheless, discrete decision functions, lacking gradient information, often require specialized training methods.
The trend in Figure 3 indicates classical statistical classification methods can address discrete decision-making problems. We can connect block features to decision-making via a statistical classifier. By classifying general features (i.e., “gap” and “top prob”), we simplify decision-making into binary classification, enabling an early exit strategy. If the classifier considers the current layer’s features stoppable, subsequent layers’ computations can be discarded; otherwise, continue to the final layer. This process is also illustrated in Figure 2(a).
4.3 Classifier Objective
Here we detail the training process of the classifier through their objectives, respectively. Given one instance, we calculate the feature vector using the feature selection module. This feature vector serves as the input for the classifier module. If the current layer’s output provides the correct answer , the associated label is a positive example; otherwise, it’s a negative example.
Thus, for an layer LLM, each input instance yields pairs of . The details of creating training data for classifier are in Appendix B. We consider two types of classifiers, Support Vector Machines (SVM) Hearst et al. (1998) and Conditional Random Fields (CRF) Lafferty et al. (2001). The first one does not rely on the context of sequences, while the second one takes into account that the features of layer-by-layer blocks might implicitly incorporate concepts of sequence modeling.
SVM Objective.
SVM aims to find an optimal hyperplane that separates classes by minimizing classification errors and maximizing the margin between support vectors.
CRF Objective.
CRF is used to capture sequence feature dependencies and make decisions based on neighboring element states in sequence labeling tasks, with the training objective of maximizing the conditional likelihood of the true label sequence given the input sequence.
@@ 5 Experiments
We now conduct experiments with AdaInfer on well-known LLMs across various tasks.
5.1 Evaluation Tasks
To evaluate the zero/few-shot learning capabilities of AdaInfer, we utilize three primary types of tasks.
@@ 5.1 Evaluation Tasks
Rule Understanding Task.
GPT-3’s Brown et al. (2020) few-shot learning capability is tested with tasks requiring pattern recognition, using synthetic datasets from Todd et al. (2024); Hernandez et al. (2024) for tasks like Capitalize/Lowercase Letter, Choose Item/Category from List, and recognizing data pairs (e.g.,, Landmark-Country).
@@ 5.2 Experiment Settings
| Setting | Model | MMLU | CommonsenseQA | SQuAD | Avg | ||||
| Acc | FLOPs | Acc | FLOPs | Acc | FLOPs | Acc | FLOPs | ||
| Zero-shot | OPT-13B | 7.95 | 100 | 8.20 | 100 | 20.00 | 100 | 12.05 | 100 |
| AdaInfer | 8.67 | 97.55 | 2.80 | 97.55 | 23.00 | 97.55 | 11.49 | 97.55 | |
| Few-shot | OPT-13B | 23.60 | 100 | 21.45 | 100 | 26.12 | 100 | 23.72 | 100 |
| AdaInfer | 22.59 | 83.94 | 21.62 | 86.05 | 25.95 | 88.31 | 23.39 | 86.10 | |
| Zero-shot | Llama2-13B | 2.54 | 100 | 1.00 | 100 | 19.20 | 100 | 7.58 | 100 |
| AdaInfer | 2.48 | 98.14 | 0.70 | 98.37 | 25.90 | 85.34 | 9.69 | 93.95 | |
| Few-shot | Llama2-13B | 53.31 | 100 | 64.92 | 100 | 52.9 | 100 | 57.04 | 100 |
| AdaInfer | 52.44 | 93.55 | 62.48 | 89.10 | 48.35 | 80.66 | 54.42 | 87.77 | |
| Setting | Model | Sentiment | AG News | Avg | Rule Understanding | ||||
| Acc | FLOPs | Acc | FLOPs | Acc | FLOPs | Acc | FLOPs | ||
| Zero-shot | OPT-13B | 0.00 | 100 | 0.10 | 100 | 0.05 | 100 | 3.38 | 100 |
| AdaInfer | 0.00 | 96.87 | 0.10 | 100 | 0.05 | 98.44 | 3.86 | 92.52 | |
| Few-shot | OPT-13B | 92.58 | 100 | 72.83 | 100 | 82.71 | 100 | 58.48 | 100 |
| AdaInfer | 92.97 | 80.28 | 72.83 | 100 | 82.90 | 90.14 | 52.83 | 89.74 | |
| Zero-shot | Llama2-13B | 0.00 | 100 | 0.10 | 100 | 0.05 | 100 | 2.32 | 100 |
| AdaInfer | 0.00 | 97.43 | 0.10 | 88.37 | 0.05 | 92.90 | 6.14 | 85.76 | |
| Few-shot | Llama2-13B | 95.90 | 100 | 77.53 | 100 | 86.72 | 100 | 69.36 | 100 |
| AdaInfer | 92.65 | 59.70 | 76.43 | 87.69 | 84.54 | 73.70 | 61.87 | 80.61 | |
@@
For AdaInfer’s backbone, we choose widely recognized LLMs, detailed in Table 1. These models vary in terms of the number of parameters, ranging from 7 billion to 70 billion, and the number of layers, ranging from 32 layers to 80 layers. Specifically, our selections encompass OPT Zhang et al. (2022) and the Llama 2 series Touvron et al. (2023). These models exhibit subtle differences in architectural design and training data volume.
@@
We evaluate our approach under zero-shot and few-shot scenarios, using sample sizes of 5, 10, 15, and 20. For zero-shot, the input is the test set’s . For few-shot, training set examples are added to . For in-context learning prompts, we use a default template: , concatenating random and samples from task-specific training sets.
@@
For performance evaluation, we report the top-1 accuracy score on the test set following Todd et al. (2024). When tokenizing into multiple tokens, we treat the first token of as the target token. To evaluate computational efficiency, we determine the early exit layer index for each input instance, which can be translated into floating-point operations (FLOPs) ratios for comparison using the method described in Narayanan et al. (2021). The FLOPs ratio is calculated as:
where represents the stop layer index during inference in AdaInfer, is the total number of transformer layers, denotes the hidden size, is the sequence length, and stands for vocabulary size. Although AdaInfer converts hidden states to logits at each block through a classification layer, it only utilizes the last token’s hidden state even with longer sequences. Hence, this computation can be ignored (0.03% of the total FLOPs for transformer inference). Further computation details on the calculation process can be found in Appendix A. Statistical classifiers have significantly lower computational costs compared to LLM inference, as detailed in Appendix A, allowing us to overlook this aspect in our analysis.
5.3 Comparison with Baseline Methods
@@
The main experimental results of AdaInfer are presented in Tables 2 and 3. These experiments were conducted in zero-shot and few-shot settings, showcasing the Top-1 accuracy and average FLOPs ratios (compared to the baseline). From a perspective of performance and computational efficiency, we can draw the following experimental conclusions.
Performance is Comparable with Minimal Loss.
@@
Tables 2 and 3 show that across both zero-shot and few-shot settings, top-1 average accuracy remains within a narrow margin of 5% for all tasks and 1% for QA and text classification task groups when compared to baseline models. AdaInfer maintains LLMs capabilities and in-context learning abilities without modifying model parameters. This finding is promising, especially in light of our observation1 in Section 3.2, where we demonstrate the feasibility of implementing early exit strategies within LLM middle layers while preserving performance. For certain tasks, AdaInfer surpasses the last layer (baseline) in zero-shot or few-shot accuracy. This hints at a tendency for deep layers to potentially over-represent certain tasks, which could impede performance during LLM inference.
Reducing FLOPs Cost from 2% to 41%.
We convert the average and variance of early exit layers for each task to FLOPs ratios in Table 2 and Table 3. It can be observed that the FLOPs ratios vary for different types of tasks, ranging from to . This variation is because AdaInfer assesses different early exit layer configurations for different task inputs. Even for the same task with different inputs, AdaInfer may recommend different early exit layer settings. For instance, in the sentiment analysis task, a reduction in computational cost can be achieved using Llama2-13B, while for the knowledge-based question answering MMLU and Commonsense question answering CommonSenseQA, the savings range from to . This aligns with our observation2 outlined in Section 3.2, where we argue that at LLM inference scenario, Not all Layers are Necessary, and allocating fewer computational resources for “simple” samples can improve computational efficiency.
@@
| Rule Understanding | Zero-shot | 9.90 | 74.80 | 3.43 | 90.29 |
| Few-shot | 53.78 | 70.38 | 65.82 | 90.29 | |
5.4 Evaluation on Different Exit Strategy
In the main experiments Table 2 and Table 3, we employ SVM as the classifier for AdaInfer. To explore the impact of different classification strategies, Table 4 compares the effects of implementing an early-exit strategy with a GAP threshold set at 0.8 (stopping computation when the current block’s GAP feature exceeds 0.8) against using CRF as a classifier. The results indicate that both GAP and CRF can reduce computational costs from 3% to 50% and maintain comparable LLM performance. Notably, in the zero-shot setting, GAP outperforms CRF, suggesting a relatively weak dependency between block features.
@@ 5.5 Evaluation across Scaling Law
In our main experiments, we employ 13B-sized Llama2 and OPT models. To explore the effects of AdaInfer on models of different sizes, we conduct experiments on the 7B and 70B versions of Llama2. The results for the 7B model, presented in Table 5, show that AdaInfer either maintains accuracy with minimal (<1%) loss or exceeds the baseline in certain tasks, and achieves a computational reduction ranging from 4% to 32%. However, in experiments with the 70B model, we observe that in a zero-shot setting, AdaInfer matches or slightly exceeds the baseline model while reducing computational costs by 10% to 50%. However, in the few-shot setting, despite similar reductions in computation, AdaInfer’s accuracy shows a 1% to 25% drop across different tasks compared to the baseline. This suggests that for larger models, such as the 70B or even larger scales, AdaInfer may need to more precisely identify and utilize features at different levels. Improving AdaInfer to adapt to these larger models is a direction for our future research. The results of all LLMs using different classifiers are summarized in Table 8 and Table 9 in the Appendix D and we have highlighted the best results for each task in the current setting.
| Task | Setting | Llama2-7B | AdaInfer | ||
| Acc | FLOPs | Acc | FLOPs | ||
| MMLU | Zero-shot | 4.19 | 100 | 4.63 | 96.13 |
| Few-shot | 43.05 | 100 | 43.73 | 93.76 | |
| CommonsenseQA | Zero-shot | 5.30 | 100 | 4.80 | 95.26 |
| Few-shot | 53.50 | 100 | 53.00 | 90.46 | |
| SQuAD | Zero-shot | 20.40 | 100 | 23.80 | 89.98 |
| Few-shot | 48.08 | 100 | 45.82 | 87.06 | |
| Sentiment | Zero-shot | 0.00 | 100 | 0.00 | 96.37 |
| Few-shot | 95.20 | 100 | 95.30 | 68.05 | |
| AG News | Zero-shot | 0.10 | 100 | 0.10 | 91.36 |
| Few-shot | 79.65 | 100 | 79.72 | 94.51 | |
| Rule Understanding | Zero-shot | 5.47 | 100 | 5.32 | 91.55 |
| Few-shot | 66.80 | 100 | 66.92 | 88.41 | |
@@ 5.6 Generalization Study
In Tables 2 and 3, we randomly select 6 to 9 training datasets from the entire pool of task training sets, which altogether contain 71 sub-datasets, to train the AdaInfer classifier. Furthermore, to assess the generalization performance of the statistical classifiers, we conduct the following tests.
@@
The results are presented in Table 6. The SVM classifier exhibits satisfactory intra-task and inter-task generalization capabilities, consistent with the results presented in the main results. However, for the CRF classifier, training in an intra-task manner leads to premature termination of the LLM at very shallow layers, resulting in subpar performance. This could be attributed to insufficient feature selection, causing the CRF to overfit noise or local features in the training data. Additionally, due to variations in the logits distribution characteristics among different models, the inter-model classifier’s performance shows moderate accuracy. In conclusion, based on the results from Table 2 and Table 3 and Table 6, when using AdaInfer, we recommend utilizing SVM as the classifier.
@@ 5.7 Factor Study
In response to the features mentioned in Section 4.1, we conduct cross-validation. Given that the classifiers in the main results utilized basic features (i.e., “gap”, “top prob”), we explore the impact of features such as the cosine similarity between the current block and the previous block, which encompasses the attention values (attn), multi-layer perceptron (mlp), and hidden states. The results are presented in Table 7. The attention values have no discernible impact on the results, while other features like mlp and hidden states have an adverse effect. This result is consistent with the trend shown in Figure 3, indicating that logits can measure whether the model’s current forward progress is sufficient, while changes in other features may involve various factors.
@@ 6 Conclusion
In this paper, we first give evidence of that Not all Layers are Necessary during Inference and provide statistical evidence to support this. Then, we present AdaInfer, a simple yet effective algorithm that determines the appropriate moment to cease inference based on the input instance, thus enhancing inference efficiency and adaptability without modifying the model’s parameters. Experiment results show that AdaInfer can reduce an average of 14.8% computational resources, even up to 50% on sentiment tasks while maintaining comparable performance. More importantly, AdaInfer is compatible with other model acceleration techniques, potentially offering further improvements in inference efficiency. We argue that AdaInfer establishes a new paradigm for efficient inference besides effective existing methods.
@@
In this paper, we make the first attempt to discover that the logits of each block are critical for early-exit classifiers in LLMs, incorporating this insight as a key design choice in AdaInfer. However, since AdaInfer relies on a single forward pass, it has not yet been extended to sequential generative tasks, offering significant avenues for future research.

### v3

@@ 4.2 Classifier
The classifier determines if the signal is compelling enough to warrant an early termination of the process. There are many choices for classifiers, ranging from rule-based classifiers Huang et al. (2017); Yang et al. (2020); Wang et al. (2022) to gating functions Lin et al. (2017); Bejnordi et al. (2019). In our context, classical statistical classification methods are a good option due to their efficiency and their ability to handle simple input features (i.e., “gap” and “top prob”) for a binary classification task.
Given one instance, we obtain the feature vector using the feature selection module. If the current layer’s output provides the correct answer , the associated label is a positive example; otherwise, it is a negative example. For LLMs trained to predict the next token, if the next token predicted based on an intermediate decoding layer’s output is the same as the token predicted by the last decoding layer’s output, then the layer’s label .
Thus, for an -layer LLM, each input instance yields pairs of . The details of creating training data for the classifier are provided in Appendix B. In our implementation, we consider two types of classifiers: Support Vector Machines (SVM) Hearst et al. (1998) and Conditional Random Fields (CRF) Lafferty et al. (2001). SVM does not rely on the context of sequences, while CRF incorporates sequence modeling along the decoding layers.
@@ 5 Experiments
We now conduct experiments with AdaInfer on well-known LLMs across various tasks. Specifically, we evaluate the zero/few-shot learning capabilities, with two primary types of tasks.
@@ 5.1 Experiment Settings
| Tasks | P. Ratio(%) | MMLU | CommonsenseQA | SQuAD | Sentiment | AG News | ||||||||||
| Acc | #Avg. L | Var | Acc | #Avg. L | Var | Acc | #Avg. L | Var | Acc | #Avg. L | Var | Acc | #Avg. L | Var | ||
| Llama 7B (32 Layers) | ||||||||||||||||
| Dense | – | 43.05 | 32 | – | 53.50 | 32 | – | 48.08 | 32 | – | 95.20 | 32 | – | 79.65 | 32 | – |
| ShortGPTp | 28.13 | 21.52 | 23 | – | 33.52 | 23 | – | 10.60 | 23 | – | 93.48 | 23 | – | 56.90 | 23 | – |
| ShortGPT5 | 15.63 | 29.95 | 27 | – | 41.90 | 27 | – | 12.97 | 27 | – | 90.40 | 27 | – | 53.25 | 27 | – |
| ShortGPT3 | 9.38 | 37.39 | 29 | – | 53.22 | 29 | – | 14.32 | 29 | – | 94.17 | 29 | – | 71.28 | 29 | – |
| AdaInfer | 9.66 35.71 | 43.73 | 28.91 | 4.97 | 53.00 | 27.90 | 5.93 | 45.82 | 26.77 | 11.88 | 95.30 | 20.57 | 5.10 | 79.72 | 29.20 | 2.70 |
| Llama 13B (40 Layers) | ||||||||||||||||
| Dense | – | 53.31 | 40 | – | 64.92 | 40 | – | 52.90 | 40 | – | 95.90 | 40 | – | 77.53 | 40 | – |
| ShortGPTp | 25 | 45.12 | 30 | – | 65.00 | 30 | – | 13.32 | 30 | – | 84.38 | 30 | – | 55.90 | 30 | – |
| ShortGPT5 | 12.50 | 46.64 | 35 | – | 64.45 | 35 | – | 16.35 | 35 | – | 89.80 | 35 | – | 70.17 | 35 | – |
| ShortGPT3 | 7.50 | 47.22 | 37 | – | 64.47 | 37 | – | 17.25 | 37 | – | 95.90 | 37 | – | 75.47 | 37 | – |
| AdaInfer | 9.13 43.33 | 52.44 | 36.35 | 8.15 | 62.48 | 34.60 | 10.20 | 48.35 | 31.18 | 31.75 | 92.65 | 22.67 | 8.10 | 76.43 | 34.02 | 24.18 |
| OPT 13B (40 Layers) | ||||||||||||||||
| Dense | – | 23.60 | 40 | – | 21.45 | 40 | – | 26.12 | 40 | – | 92.58 | 40 | – | 72.83 | 40 | – |
| ShortGPTp | 25 | 10.17 | 30 | – | 11.50 | 30 | – | 0.65 | 30 | – | 14.72 | 30 | – | 2.27 | 30 | – |
| ShortGPT5 | 12.50 | 22.92 | 35 | – | 19.12 | 35 | – | 22.12 | 35 | – | 86.33 | 35 | – | 49.42 | 35 | – |
| ShortGPT3 | 7.50 | 23.05 | 37 | – | 19.68 | 37 | – | 24.65 | 37 | – | 91.35 | 37 | – | 66.62 | 37 | – |
| AdaInfer | 9.75 22.63 | 22.59 | 32.37 | 7.92 | 21.62 | 33.33 | 12.12 | 25.95 | 34.20 | 13.50 | 92.97 | 30.95 | 5.77 | 72.83 | 39.00 | 0.00 |
@@
We evaluate AdaInfer under zero-shot and few-shot scenarios, using sample sizes of 5, 10, 15, and 20. For zero-shot, the input is the test set’s . For few-shot, training set examples are added to . For in-context learning prompts, we use a default template: , concatenating random and samples from task-specific training sets.
@@
We report the top-1 accuracy score on the test set following function vectors Todd et al. (2024) (HELM implementation)111https://huggingface.co/blog/open-llm-leaderboard-mmlu. For computational efficiency, we follow previous work Ma et al. (2023); Schuster et al. (2022); Elbayad et al. (2019) and report the pruning ratio (P. Ratio) and the average number of activated layers (#Avg. L) for each task, along with their variance (Var). These metrics directly measure complexity reduction, avoiding conflation with implementation or infrastructure-specific details Dehghani et al. (2021). For reference, we also translate them into FLOPs reduction ratios, in the Appendix D. Considering the conditional checks and classifier computation involved in AdaInfer, we also compare the actual speed of AdaInfer in real-world scenarios with Dense implementation, reporting wall-clock time Dehghani et al. (2021).
Baseline Method: ShortGPT.
We compare AdaInfer with the structured pruning method ShortGPT Men et al. (2024), which prunes redundant layers in LLMs based on similarity scores. For the OPT model, we calculate redundant layers as outlined in the paper. For the LLama model, we use the same layers reported. Note that these model pruning methods apply a static pruning ratio across all tasks, whereas our AdaInfer adaptively performs model pruning based on input.
5.2 Main Results
@@
The main results of AdaInfer are presented in Table 2. Conducted in few-shot settings, these experiments show the Top-1 accuracy, pruning ratios, average active layers for each task, and their variance. From the perspective of performance and computational efficiency, we draw the following experimental conclusions.
AdaInfer has minimum impact on performance (<1%).
Table 2 shows that the Top-1 accuracy of AdaInfer remains within a very narrow margin of less than 1% for all tasks compared to dense models, i.e., without early exit. In contrast, ShortGPT, following the paper’s setting and denoted as ShortGPTp, experiences a significant performance drop for almost all tasks 222We noted a decline in the performance of the reproduced ShortGPT on the SQuAD dataset when the prompts increased to 10, 15, 20 shots.. Since AdaInfer adaptively skips decoding layers, the number of layers skipped varies for different instances and across different tasks. For a fair comparison, we have also evaluated ShortGPT5 and ShortGPT3, which skip the last 5 and 3 decoding layers, respectively. The numbers of skipped layers are chosen to match the overall range of layers skipped by AdaInfer. This allows for a more comprehensive comparison with methods that use a fixed pruning ratio Yang et al. (2024); Ma et al. (2023); Men et al. (2024). The results in Table 2 demonstrate that AdaInfer surpasses both settings.
@@
In short, AdaInfer achieves adaptive inference while maintaining LLM capabilities and in-context learning abilities without modifying model parameters. This finding is promising, especially in light of our Observation 1, where we demonstrate the feasibility of implementing early exit strategies while preserving performance. As shown in Table 2, AdaInfer even surpasses the last layer accuracy for certain tasks. This suggests that deep layers may over-represent certain instances, potentially impeding performance during LLM inference.
Pruning ratio ranges from 9% to 43%, average 17.8%.
We report the average and variance of the activated layers for each task and compute the pruning ratios in Table 2. The pruning ratios vary from task to task, ranging from to , a clear indication of AdaInfer assesseing different early exit layer configurations for different task inputs. More layers are skipped for simple tasks like sentiment analysis task, where a reduction in computational cost can be achieved on Llama2-13B. For more complex question answering tasks, the savings range from to .
Wall-clock time.
| Task | Llama2 7B (FP32) | Llama2 13B (FP16) | ||||
| Dense | Ours | Speed up | Dense | Ours | Speed up | |
| MMLU | 796.53 | 781.31 | 1.02x | 339.19 | 320.46 | 1.05x |
| Sentiment | 41.18 | 39.69 | 1.04x | 28.18 | 21.76 | 1.30x |
Next, we study the end-to-end runtime of AdaInfer. Table 3 compares the runtime of AdaInfer with a dense implementation on MMLU and Sentiment tasks (5-shot, batch size set to 1), using (32GB). We observed a 1.03x speed up on MMLU and 1.17x speed up on Sentiment when applying AdaInfer. That is, despite AdaInfer converting hidden states to logits at each block through the LM head layer, it only utilizes the last token’s hidden state, which is independent of the input sequence length. Consequently, this computation cost is minimal (0.03% of the total FLOPs for transformer inference). Further computational details on this process can be found in Appendix A. Meanwhile, statistical classifiers like SVM have much lower computational costs compared to LLM inference, as detailed in Appendix A, highlighting the computational efficiency potential of AdaInfer.
@@
5.3 Evaluation on Alternative Exit Strategies
In Table 2, we employ SVM as the classifier for AdaInfer. To explore the impact of alternative exit strategies, Table 4 reports AdaInfer implemented with a GAP threshold set at 0.8 (stopping inference when the current block’s GAP feature exceeds 0.8) and AdaInfer with CRF as the classifier. The results show that both GAP and CRF can reduce computational costs by 3% to 50% while maintaining comparable LLM performance. Notably, in the zero-shot setting, GAP outperforms CRF, suggesting a relatively weak dependency between block features.
@@ 5.4 Evaluation across Scaling Law
Table 2 reports results on 7B/13B-sized Llama2 and OPT models. In experiments with the Llama2 70B version, we observe that in a zero-shot setting, AdaInfer matches or slightly exceeds the dense model while reducing computational costs by 10% to 50%. However, in the few-shot setting, despite similar reductions in computation, AdaInfer’s accuracy shows a 1% to 25% drop across different tasks compared to the dense model, i.e., without early exit. This calls for more feature engineering for larger models, such as the 70B or even larger scales. Improving AdaInfer to adapt to these larger models is a direction for our future research. The results of all LLMs using different classifiers are summarized in Table 7 and Table 8 in Appendix D. The best results are highlighted for each task in the current setting.
@@ 5.5 Generalization Study
In Tables 2, we randomly select 6 training datasets from the entire pool of task training sets, which altogether contain 71 sub-datasets, to train the AdaInfer classifier. Furthermore, to assess the generalization performance of the statistical classifiers, we conduct the following tests.
@@
The results are presented in Table 5. The SVM classifier exhibits satisfactory intra-task and inter-task generalization capabilities, consistent with the results presented in the main results. However, for the CRF classifier, training in an intra-task manner leads to premature termination of the LLM at very shallow layers, resulting in subpar performance. This could be attributed to insufficient feature selection, causing the CRF to overfit noise or local features in the training data. Additionally, due to variations in the logits distribution characteristics among different models, the inter-model classifier’s performance shows moderate accuracy. In conclusion, based on the results from Tables 2 and 5, SVM is the choice of classifier for AdaInfer.
@@ 5.6 Factor Study
In response to the features identified in Section 4.1, we conduct cross-validation. Given that the classifiers in the main results utilized basic features (i.e., “gap”, “top prob”), we explore the impact of features such as the cosine similarities between the current block and the previous block, which encompasses the attention values (attn), multi-layer perceptron (mlp), and hidden states. Results in Table 6 show that attention values have no discernible impact on the results, while other features like mlp and hidden states have an adverse effect. This result is consistent with the trend shown in Figure 3. It is our understanding that logits can measure whether the model’s current forward progress is sufficient, while changes in other features may involve various factors.
@@ 6 Conclusion
In this paper, we first provide an analysis to show that not all layers are necessary during inference for LLMs. Then, we present AdaInfer, a simple yet effective algorithm that dynamically determines the appropriate moment to cease inference based on the input instance. The decision is predicted by a low-cost statistical classifier using two easily obtainable features: the probability estimated for the most likely token and the gap between this probability and that of the next most likely token. While these two features may not capture all the evidence needed for early exit, the results are very promising for enhancing inference efficiency and adaptability, without modifying the model’s parameters. Experiments on well-known LLMs (e.g., Llama2 series and OPT) show that AdaInfer can achieve an average pruning ratio of 17.8%, and up to 43% on sentiment tasks, with nearly no performance drop (<1%). The computational savings can be more significant if an LLM is deployed to process mostly simple tasks and only occasionally difficult ones. Furthermore, AdaInfer is compatible with other model acceleration techniques, potentially offering further improvements in inference efficiency. We argue that AdaInfer establishes a new paradigm for efficient inference alongside existing effective methods.
@@
In this paper, we make the first attempt to discover that the logits of each block are critical for early-exit classifiers in LLMs, incorporating this insight as a key design choice in AdaInfer. However, since AdaInfer relies on a single forward pass, it has not yet been extended to sequential generative tasks, offering significant avenues for future research. Further, the may exisit more effective features in addition to logits.